Evening shift, 15 Sep 1999

Leo Bellantoni (bellanto@fnal.gov)
Thu, 16 Sep 1999 00:25:53 -0500 (CDT)

Shift Summary for EVE shift, 15/SEP/99

On Shift:
Leo Bellantoni, Wah

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 2.0

Accounting of lost time:

========================================================================
Runs ended this shift:
None

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift:
Many accesses for E832 -> E799 changeover. See below.

========================================================================
Overview of shift:
E.g., calibrate CsI, E832 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 Began reinflation of He bag2
15:52 Disconnected BA3 from FERA readout
16:00 AC has been back on long enough to lower SDAQ
temperature alarm to regular value. More work
will be needed tomorrow. Survey crews are gone;
they will resume tomorrow. List of progress
with surveys in logbook XXI, pg 193.
17:00 DC voltages back on
17:22 Power glitch. Tripped Vacuum pumps, FastBus
racks w/ bananas, several CsI MF, Pipeline
crate 7. All is back up by 19:25 or so...
Angela's GC run on the strange DC gas was
clobbered.
18:55 Reg HV turned off from at front panel of MF 15.
??? New 3MU3Y trigger installed and timed in
??? HA-1 and HA-2 thresholds adjusted to 3 and 6
mips, respectively.
14528 20:47 Cosmics run started for fruit debugging; we had to
pull KQ4X out of the readout at first, put it back
in later though.
21:45 Pipeline crate 7 tripped again. Also, MF 6 is
off.
14529 22:23 A short but successful startup of E799
configuration.
14530 22:50 Back to running cosmics

========================================================================
Follow up on Notes from Previous Shift

E832/E799 cake finished off at 17:22

========================================================================
Items to Follow up on Next Shift

Continue to keep an eye on KTeV hall temp; Et is still off in anticipation
of chilled water power down early tommorrow AM.

DC HV to be lowered, electronics work by Hogan on owl shift.

Continue surveys in morning, perhaps salvage GC of DC gas(?), reduce
aperture of 90 degree monitor. Reconnect & debug DC4 DCORs, latches
for DC3 & DC4. More fruit debugging, testing of E799 readout. After
survey, reassemble NM2 vacuum line, bring up TRD HV, move mask, REG out.

I think that is it!

Leo
-------------------------------------------------------------------------------
Leo Bellantoni WH13SX, FermiLab Batavia, IL 60510
Phone:(630)840-8772 Pager:(630)266-3803
"The more I use computers, the more I like pencils."
-------------------------------------------------------------------------------